12. Madison Company started a share appreciation plan on January 1, 2015, when it granted 210,000 rights to its executives. TDate December 31, 2015 December 31, 2016 Closing Market Price $ $ 61
Answer #1

The fair value of share on December 2015 is 66, in other words the share value went up by 17. (66-49)

The executive will have an expense of 17 times 210000 which is equal to 3570000.

The vesting period is 2 years so we have to split the expense for two years 3570000/2 = 1785000

The compensation expense in 2015 is 1,785,000

In December 2016 share value went down by 5, (69-66)

The current year expense will be (61-49) x 210000 = 2,520,000

Opening provision is 1785000

So provision for the last year will be 2520000-1785000 = 735000

The compensation expense in 2016 is 735,000

(note that we will make 100% Provision for the last year)

JOURNAL ENTRIES

FOR 31 DECEMBER 2015

Compensation Expense 1,785,000

Obligation under SAR Plan 1,785,000

FOR DECEMBER 2016

Compensation Expense 735,000

Obligation under SAR Plan 735,000

FOR JANUARY 2017

Obligation under SAR Plan 2,520,000

Cash     2,520,000

____________________________________________________________________

If you need any doubt or clarification give me a comment.

If you are satisfied with the answer give this answer a thumps up and let me know

THANK YOU!